OpenStudy (anonymous):

I will fan and medal!!!!!!! Renee took a long flight to visit her family that lives out of the country. She slept for the first quarter of the trip and watched movies for half of the remaining flight. If the total trip took nine and three-fourths hours, how long did Renee watch movies?

OpenStudy (anonymous):

9and3/4 hors = 585 mins 1/4 of 585 = 146.25 585-146.25 = 438.75 movies watched = 438.75/2 = 219.375 mins

OpenStudy (anonymous):

I don't know what that mean???

OpenStudy (anonymous):

Here!! Renee took a long flight to visit her family that lives out of the country. She slept for the first quarter of the trip and watched movies for half of the remaining flight. If the total trip took nine and three-fourths hours, how long did Renee watch movies? Choose one answer. A. three and twenty-one thirty-seconds hours B. three and one twenty-fourth hours C. three and five-twelfths hours D. five and two-thirds hours

OpenStudy (anonymous):

option A

OpenStudy (anonymous):

ok..thx